OpenStudy (anonymous):

Please Help! Will give medal and become fan Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y=2x - 17 y= -1/2x - 2 y=2x - 2 y=-1/2x - 7

OpenStudy (anonymous):

The equation you'll want to use here is: \[y-y_1 = m(x-x_1)\]Where m is the slope, x_1 and y_1 are the coordinates of any point on the line. Since the line is parallel to y=2x+2, it must have the same slope, so m=2. Now, just plug in the slope and your point (5,3) and simplify.

OpenStudy (anonymous):

So it will be the second one?

OpenStudy (mathstudent55):

Parallel lines have the same slope. The line you have is in the slope-intercept form, \(y = mx + b \) where \(m = \) slope, and \(b = \) y-intercept

OpenStudy (anonymous):

No, the second one has a slope of -1/2 and not 2.

OpenStudy (anonymous):

Then the first one because it has the slope of 2

OpenStudy (mathstudent55):

Now look at your line: \(y = 2x + 2\) When you compare it with \(y = mx + b\), you see that the slope m = 2. Since parallel lines have the same slope, the line you are looking for also has a slope of 2. That means the line you need has the equation: \(y = 2x + b\) Now we need to find b. Since you have a given point, \((5, 3)\), insert the x- and y-coordinates of the point into \(y = 2x + b\) and solve for b: \(y = 2x + b\) \(3 = 2(5) + b\) \(3 = 10 + b\) \( -7 = b\) Now that you know that \(b = -7\), insert that value in the equation \(y = 2x + b\) getting: \(y = 2x - 7\)

OpenStudy (anonymous):

so that wouldnt be it?

OpenStudy (mathstudent55):

Perhaps the original poster has a typo. I think choice A shoud read \(y = 2x - 7\), not \(y = 2x - 17\).

OpenStudy (anonymous):

Thank you so much

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

Can you do another problem please

OpenStudy (mathstudent55):

Yes. Please start a new post.

OpenStudy (anonymous):

ok